OpenStudy (anonymous):

How can I find the equation of a parabola with vertex (4,3), focal length of 1/2 and opens down?

OpenStudy (anonymous):

Since it opens down I know it is (x-4)^2=-...(y-3) but I don't know what to do with the focal length

OpenStudy (mertsj):

Since it opens down it is of the form x^2=4py

OpenStudy (anonymous):

Wait focal length is p right? So 2

OpenStudy (mertsj):

p = 1/2 and 4p = 2 So: x^2=2y Now apply the shift for the vertex: (x-4)^2=2(y-3)

OpenStudy (anonymous):

It opens down, so wouldn't it be (x-4)^2=-2(y-3)?

OpenStudy (mertsj):

No. -2y-3 Without the parentheses

OpenStudy (anonymous):

ty

OpenStudy (mertsj):

yw